OpenStudy (anonymous):

Solve for x (under sqrt) x+1=5

OpenStudy (anonymous):

If you square both sides you get rid of the sqrt, leaving you with x+1 = 5^2, and you just solve for x...x=24.

OpenStudy (anonymous):

x=24

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

\[\sqrt{x+1} = 5\]\[x+1 = 5^2 = 25\]\[x = 24\]